# MAAdvisor: Zero-Shot Index Advisor using Multi-Agent LLMs

### v1

@@ Abstract.
Index recommendation is one of the most important problems in database management system (DBMS) optimization. Given queries and certain index-related constraints, traditional methods rely on heuristic optimization or learning-based models to select effective indexes and improve query performance. However, heuristic optimization suffers from high computation time, and learning-based models lose generalisability due to training for different workloads and database schemas. With the recent rapid development of large language models (LLMs), methods using prompt tuning have been proposed to enhance the efficiency of index selection. However, such methods still can not achieve the state-of-the-art (SOTA) results, and preparing the index selection demonstrations is also resource-intensive. To address these issues, we propose MAAdvisor, a zero-shot LLM-based index advisor with a multi-agent framework. We decompose the index recommendation problem into sub-steps, including planning, selection, combination, revision, and reflection. A set of LLM-embedded agents is designed to handle each one of the different sub-steps. Our method utilizes global agents to control the index selection process and local agents to select and revise indexes. Through extensive experiments, we show that our proposed MAAdvisor not only achieves the SOTA performance compared to the heuristic methods, but also outperforms learning-based and prompt-based methods with higher efficiency and better zero-shot inference ability.
@@ 1. Introduction
Indexes in database systems significantly accelerate data retrieval and reduce redundant I/O operations by providing alternative access paths to data based on attribute values. With the appropriate index configuration, database systems can better support analytical workloads that are highly demanded by many modern applications, such as inventory management, online shopping, and real-time advertising. However, selecting and creating indexes always comes with significant computation time and storage cost. Therefore, given the constraints such as limited resource and budget, it is important for database administrators (DBAs) to select the optimal indexes that can meet the data management needs.
| Heuristic-based method | Learning-based method | Prompt-based method | MAAdvisor (Ours) | |
@@
To assist DBAs with index selection, index advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Lan et al., 2020; Perera et al., 2021; Sadri et al., 2020a, b; Skelley, 2000; Wang et al., 2024b; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997; Kossmann et al., 2022; Zhao et al., 2025; Zhou et al., 2022; Paludo Licks et al., 2020) are often adopted to automatically select the optimal indexes for a given query workload on the target database. Existing index advisors fall into three main categories: heuristic-based advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Skelley, 2000; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997), learning-based advisors (Lan et al., 2020; Perera et al., 2021; Sadri et al., 2020a, b; Wang et al., 2024b; Kossmann et al., 2022; Zhou et al., 2022; Paludo Licks et al., 2020), and prompt-based LLM advisors (Zhao et al., 2025). Heuristic-based advisors typically use rule-based or greedy algorithms to identify optimal indexes based on database optimizers’ cost estimations. However, given the large number of columns involved in complex workloads, exploring exhaustively the search space for all possible index combinations becomes prohibitory expensive. In addition, the reliance on heuristics and optimizers’ estimations can lead to suboptimal index selection. As highlighted in (Ding et al., 2019), the errors in cost estimations can even cause workload performance degradation, a phenomenon known as the index regression problem. Learning-based advisors have emerged as a promising alternative to traditional heuristic-based approaches. They utilize techniques like deep learning and reinforcement learning to train index selection models and index benefit estimators. Through extensive learning on features extracted from queries, execution plans, and database statistics, these models can capture complex interactions between indexes and queries that are difficult to encode manually and hence achieve SOTA performance. However, training these models often requires a large amount of labeled data and time. The lack of ability to generalize to unseen database limits their applicability in dynamic or evolving environments. Moreover, the need for data annotations and model re-training poses a major limitation in data-sensitive domains like healthcare or security, where data collection and sharing are tightly regulated. To increase advisors’ generalisation and reduce recommendation time cost, prompt tuning on LLMs is adopted (Zhou et al., 2023). Such a method prompts LLMs with preprocessed index recommendation demonstrations and lets the LLM select columns to build the indexes. Although the generalisation and efficiency of advisors benefit from LLM, there are some limitations. First, the performance of the model is highly sensitive to the prompt design, such as the choice and format of demonstrations. Second, demonstrations on the analyzed database schema are required to achieve optimal results, therefore, preparing high-quality demonstrations is time-consuming and often requires expert domain knowledge. Lastly, inference with LLMs can be computationally expensive for large-scale workloads. Table 1 summarizes the characteristics and drawbacks of three types of advisors as discussed above.
@@
Given the limitations of existing methods, we aim to design an index advisor that addresses the following fundamental challenges in index recommendation: (1) Zero-shot Generalization: To accommodate the diverse landscape of real-world database deployments, an effective index advisor must exhibit strong generalization capabilities. Specifically, the ability to recommend optimal indexes for previously unseen databases and workloads without requiring retraining or fine-tuning on domain-specific data. (2) Effectiveness: The advisor should deliver SOTA recommendation quality, producing index configurations that demonstrably improve query performance compared to existing approaches. (3) Efficiency: The advisor must exhibit low computational overhead, enabling rapid index recommendations that scale to large workloads without imposing significant time costs.
To address the challenges above, we propose our multi-agent index advisor based on LLM, named MAAdvisor. First, to address the zero-shot generalization challenge, we build upon the success of recent multi-agent work solving other problems under the zero-shot setting (Liu et al., 2025; Li et al., 2025; Jha et al., 2025). Multi-agent frameworks can enhance scalability by dividing complex problems into manageable components processed in parallel by specialized zero-shot agents. To address the difficulty of decomposing the index recommendation problem into sub-problems, we innovatively design a hierarchical multi-agent pipeline including global agents and local agents. Global agents decompose the problem into iterative sub-steps and plan the use of local agents in each sub-step. Local agents will then perform three key functions: (1) select candidate index, (2) revise the current index selection, and (3) construct composite indexes. This design exploits LLMs’ demonstrated reasoning capabilities in index recommendation and enables the system to generalize across diverse database schemas and workload patterns in zero-shot. Second, to solve the effectiveness challenge caused by the index regression problem in most advisors, we design two novel components for our index advisor: (1) a new workload representation paradigm for the local selection agent to accurately select optimal indexes, and (2) a dedicated local revision agent that identifies and eliminates potentially regressive index recommendations by analyzing estimation errors and performance trade-offs. Third, to guarantee the computational efficiency, we constrain the search space through a budget-aware planning mechanism where the global planning agent uses storage constraints as an optimization objective, effectively pruning the solution space explored by local agents.
@@
We conduct extensive experiments and recommend indexes for multiple popular datasets and workloads, including TPC-H, TPC-DS, DSB, and JOB. We observe that MAAdvisor’s benefit-to-cost ratio outperforms all the heuristic-based baselines, which are also zero-shot, in 10 out of the 16 settings. In addition, MAAdvisor significantly outperforms the learning-based methods, which require intensive training. Compared with the few-shot prompt-based baseline, MAAdvisor also achieves better performance with lower inference time and no domain-specific data.
@@
To the best of our knowledge, we are the first to propose MAAdvisor, a zero-shot index advisor using LLMs, achieving SOTA results with higher efficiency and zero-shot generalization.
@@
We design a multi-agent framework to solve the index recommendation problem. The problem is decomposed into multiple sub-steps, which are much easier to solve by a cooperation of the global and local agents.
@@ 2. Preliminary
In this section, we will first clarify the concepts related to index recommendation in Section 2.1. Next, we will then explain the problem of an index advisor in Section 2.2.
2.1. Workload, Index Candidate and Budget
Workload: The workload is referred to as a set of queries that will be executed in the database. The ultimate goal of the index recommendation problem is to optimize the execution efficiency of the entire workload instead of any specific query within it.
Index Candidate: Most of the methods that solve the index recommendation problem start with collecting a set of index candidates to select from. Each candidate within this set usually stands for a table column queried by any of the queries in the workload.
Storage Budget: Index selection in databases is subject to various practical constraints, most notably storage budget. Creating indexes consumes additional disk space and incurs extra cost during data modifications. During index selection, it is significant to balance performance gains against the increased resource usage, which necessitates prioritizing indexes that offer the highest benefit-to-cost ratio.
@@ 2.2. Index Advisor
Index advisors receive the workload of a specified database along with constraints like the maximum number of allowable indexes and storage budget, and then automatically generate an index recommendation. The functionality of an index advisor can be formally defined as follows:
@@
(Index Advisor): Given a database , a workload consisting of queries querying the database and a storage budget , an index advisor will automatically compute the column candidates and construct a set of indexes with columns , optimizing the cost of the workload by
where calculates the execution time of the workload querying the database with the indexes created, and calculates the total storage cost to construct the indexes selected.
@@ 3. Related Work
3.1. Index Advisors
@@
Heuristic-based Index Advisors: Heuristic-based advisors iteratively and greedily select candidate indexes until the budget constraint is met. Within each optimization step, the index advisors will update the index recommendation according to the estimated benefits of the index candidates. There are mainly two heuristics for the index advisors to follow: the ADD heuristic and the DROP heuristic. As an example of the ADD heuristic, Extend (Schlosser et al., 2019) starts with no indexes and adds one index at a time according to the highest benefit-to-cost ratio. In contrast, Drop (Choenni et al., 1993), following the DROP heuristic, starts with all index candidates and iteratively removes the least beneficial one until optimization constraints are met. In general, heuristic-based advisors generally take significantly longer computation time to recommend indexes, especially for databases with complicated schemas. In addition, heuristic-based methods rely solely on database optimizers to evaluate index benefits. According to (Ding et al., 2019), wrong estimations in column cardinality may mislead index advisors to select indexes that are not beneficial to the workload.
Learning-based Index Advisors: While recent work (Zhou et al., 2024) shows that supervised learning approaches such as QueryFormer (Zhao et al., 2022) can be applied to index recommendation, state-of-the-art learning-based advisors (Kossmann et al., 2022; Wang et al., 2024b; Lan et al., 2020; Sadri et al., 2020a; Paludo Licks et al., 2020) predominantly adopt reinforcement learning (RL) algorithms. For example, DQN (Lan et al., 2020) integrates Q-learning with deep neural networks to derive optimal policies from workload representations, while SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b) leverage the PPO (Schulman et al., 2017) algorithm for policy optimization, with BALANCE further employing a transfer learning framework to exploit historical workloads. RL-based advisors typically achieve much lower inference latency than heuristic-based counterparts; however, the training time for such methods cannot be overlooked. In addition, the generalisability of the RL-based methods is lower since training is required for each unseen database.
Prompt-based Index Advisors: Large Language Models (LLMs) have recently emerged as a hot topic in machine learning research. These models have demonstrated a surprisingly strong ability to handle a variety of text-related tasks such as generation, decision-making, and reasoning. Therefore, (Zhao et al., 2025) proposes to utilize LLMs for index recommendation. They construct and offer in-context demonstrations for LLMs to learn from. LLMs will then directly recommend indexes for both met and unmet databases based on the demonstrations. When a new workload is given, LLMIdxAdvis explores the prepared demonstration pool and selects the optimal demonstration for the workload by cosine similarity. LLMIdxAdvis successfully addresses the problem of efficiency and robustness in index recommendation. However, the efficiency and generalisability are achieved at the cost of the effectiveness of the indexes recommended. LLMIdxAdvis achieves only comparable performance against other methods in simpler workloads and underperforms the SOTA methods given more complicated workloads and database schemas.
3.2. Multi-agent Pipelines
Leveraging the strong generalizing and reasoning abilities of LLMs, recent research proposes the idea of agents (Shen et al., 2023; He et al., 2016). LLM agents employ a more complex, modular architecture, enabling end-to-end problem-solving and the ability to handle multi-step tasks. In a standard multi-agent pipeline (Liu et al., 2025; Li et al., 2025; Jha et al., 2025), a complex problem is divided into different sub-problems. The sub-problems are assigned to specialized agents that collaborate in parallel. Each agent will focus on one task, such as data pre-processing, information retrieval, reasoning, or generation. Communication occurs through shared memory or message passing among different agents. The agents can be either rule-based or learning-based, which improves scalability and interpretability, enabling efficient problem-solving for complex workflows requiring coordination, specialization, and iterative refinement among agents.
4. MAAdvisor Pipeline Overview
Figure 1 illustrates the pipeline of MAAdvisor. In this section, we will explain our motivations in the method design and offer an overview of the two key sections in our method, named workload representation (4.1) and multi-agent pipeline (4.2).

### v2

# AMAZe: A Multi-Agent Zero-shot Index Advisor for Relational Databases

@@ Abstract.
Index recommendation is one of the most important problems in database management system (DBMS) optimization. Given queries and certain index-related constraints, traditional methods rely on heuristic optimization or learning-based models to select effective indexes and improve query performance. However, heuristic optimization suffers from high computation time, and learning-based models lose generalisability due to training for different workloads and database schemas. With the recent rapid development of large language models (LLMs), methods using prompt tuning have been proposed to enhance the efficiency of index selection. However, such methods still can not achieve the state-of-the-art (SOTA) results, and preparing the index selection demonstrations is also resource-intensive. To address these issues, we propose AMAZe, a zero-shot LLM-based index advisor with a multi-agent framework. We decompose the index recommendation problem into sub-steps, including planning, selection, combination, revision, and reflection. A set of LLM-embedded agents is designed to handle each one of the different sub-steps. Our method utilizes high-level agents to control the index selection process and low-level agents to select and revise indexes. Through extensive experiments, we show that our proposed AMAZe not only achieves the SOTA performance compared to the heuristic methods, but also outperforms learning-based and prompt-based methods with higher efficiency and better zero-shot inference ability.
@@ 1. Introduction
Indexes in database systems significantly accelerate data retrieval and reduce redundant I/O operations through providing alternative access paths to data based on attribute values. With the appropriate index configuration, database systems can better support analytical workloads that are highly demanded by many modern applications such as inventory management, online shopping and real-time advertisement. However, selecting and creating indexes always comes with significant computation time and storage cost. Therefore, given the constraints such as limited resource and budget, it is important for database administrators (DBAs) to select the optimal indexes that can meet the data management needs.
| Heuristic-based method | Learning-based method | Prompt-based method | AMAZe (Ours) | |
@@
To assist DBAs with index selection, index advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Lan et al., 2020; Perera et al., 2021a; Sadri et al., 2020a, b; Skelley, 2000; Wang et al., 2024b; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997; Kossmann et al., 2022; Zhao et al., 2025; Zhou et al., 2022; Paludo Licks et al., 2020; Perera et al., 2021b; Wu et al., 2022; Wang and Yoneki, 2024; Chang et al., 2024) are often adopted to automatically select the optimal indexes for a given query workload in the target database. Existing index advisors fall into three main categories: heuristic-based advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Skelley, 2000; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997), learning-based advisors (Lan et al., 2020; Perera et al., 2021a; Sadri et al., 2020a, b; Wang et al., 2024b; Kossmann et al., 2022; Zhou et al., 2022; Paludo Licks et al., 2020; Perera et al., 2021b; Wu et al., 2022; Wang and Yoneki, 2024; Chang et al., 2024), and prompt-based LLM advisors (Zhao et al., 2025). Heuristic-based advisors typically use rule-based or greedy algorithms to identify optimal indexes based on database optimizers’ cost estimations. However, given the large number of columns involved in complex workloads, exploring exhaustively the search space for all possible index combinations becomes prohibitory expensive. In addition, the reliance on heuristics and optimizers’ estimations can lead to suboptimal index selection. As highlighted in (Ding et al., 2019), the errors in cost estimations can even cause workload performance degradation, a phenomenon known as the index regression problem. Learning-based advisors have emerged as a promising alternative to traditional heuristic-based approaches. They utilize techniques like deep learning and reinforcement learning to train index selection models and index benefit estimators. Through extensive learning on features extracted from queries, execution plans, and database statistics, these models can capture complex interactions between indexes and queries that are difficult to encode manually and hence achieve SOTA performance. However, training these models often requires a large amount of labeled data and time. The lack of ability to generalize to unseen database limits their applicability in dynamic or evolving environments. Moreover, the need for data annotations and model re-training poses a major limitation in data-sensitive domains like healthcare or security, where data collection and sharing are tightly regulated. To exploit the adaptability and semantic reasoning strengths of LLMs while mitigating both database optimizer dependency and data dependency, prompt tuning on LLMs (Zhou et al., 2023) is adopted. Such a method (Zhao et al., 2025) prompts LLMs with preprocessed index recommendation demonstrations and allows the LLM select columns to build the indexes. While this LLM-based approach enhances the generalization and efficiency of index advisors, it also presents notable limitations. Actually, prompt tuning approach exhibits limited effectiveness in zero-shot setting due to LLMs’ limited understanding of database-specific contexts and optimization nuances without concrete examples, necessitating few-shot learning with domain-specific demonstrations (Brown et al., 2020). However, preparing high-quality demonstrations is challenging and time-consuming. In addition, inference with LLMs can be computationally expensive for large-scale workloads. When processing complex workloads, the number of tokens to be encoded and reasoned over grows rapidly, leading to high memory consumption and long inference times. This computational overhead becomes particularly pronounced in online environments. Table 1 summarizes the characteristics and drawbacks of three types of advisors as discussed above.
@@
Given the limitations of existing methods, we aim to design an index advisor that achieves the following key objectives in index recommendation: (1) Zero-shot Generalization: To accommodate the diverse landscape of real-world database deployments, an effective index advisor must exhibit strong generalization capabilities. Specifically, the ability to recommend optimal indexes for previously unseen databases and workloads without requiring retraining or fine-tuning on domain-specific data. (2) Effectiveness: The advisor should deliver SOTA recommendation quality, producing index configurations that demonstrably improve query performance compared to existing approaches. (3) Efficiency: The advisor must exhibit low computational overhead, enabling rapid index recommendations that scale to large workloads without imposing significant time costs. This efficiency is critical because database systems often handle massive datasets and complex query patterns where slow workload processing and index recommendation can delay essential optimizations, potentially impacting overall system performance and user experience.
In order to achieve the objectives above, we propose AMAZe, a multi-agent zero-shot index advisor based on LLM. To equip AMAZe with zero-shot inference capability, we build upon recent advances in multi-agent frameworks that have demonstrated success in addressing other tasks under the zero-shot setting (Liu et al., 2025; Li et al., 2025; Jha et al., 2025). A multi-agent framework can better support zero-shot inference because it decomposes the complex index selection problem into specialized sub-tasks, allowing different agents to collaborate and compensate for the limitations of a single LLM. These sub-tasks enable the framework to leverage the general knowledge encoded in LLMs while structuring their reasoning in a way that aligns more closely with database-specific requirements. We further design three critical components within the multi-agent framework: efficient workload representation, budget-aware agent planning, and effectiveness-aware agent coordination. The efficient workload representation facilitates streamlined communication among agents and mitigates LLM hallucination by providing a concise yet informative summary of the workload. Complementarily, budget-aware agent planning enhances overall efficiency by guiding the LLM to select indexes that optimize resource usage while respecting storage and computation constraints. To tackle the index regression problem, which constitutes the primary barrier to achieving effectiveness, we design effectiveness-aware agent coordination. Through cooperative interactions, multiple agents iteratively refine the current index selection by evaluating index benefits, thereby ensuring a more robust, stable, and effective index recommendation pipeline.
@@
We conduct extensive experiments and recommend indexes for multiple popular datasets and workloads, including TPC-H, TPC-DS, DSB, and JOB. We observe that AMAZe’s benefit-to-cost ratio outperforms all the heuristic-based baselines, which are also zero-shot, in 10 out of the 16 settings, and matches the best of them in the remaining 6 settings. In addition, AMAZe significantly outperforms the learning-based methods which requires intensive training. Compared with the few-shots prompt-based baseline, AMAZe also achieves better performance with lower inference time without relying on domain-specific data.
@@
We propose AMAZe, which achieves SOTA results with higher efficiency and zero-shot generalization. To the best of our knowledge, AMAZe is the first LLM-based zero-shot index advisor.
@@
We design a multi-agent framework to solve the index recommendation problem. The problem is decomposed into multiple sub-steps, which are much easier to solve by a cooperation of the high-level and low-level agents.
Extensive experiments on various workloads demonstrate that AMAZe achieves superior performance with lower inference time compared to baseline methods.
@@ 2. Preliminary
In this section, we will first clarify the concepts related to index recommendation in Section 2.1. Next, we will then formally define the problem of an index advisor in Section 2.2.
2.1. Workload, Index Candidate and Constraints
Workload: The workload is referred to as a set of queries that will be optimized in the database. The ultimate goal of the index recommendation problem is to optimize the execution efficiency of the entire workload instead of any specific query within it.
Index Candidate: Most of the methods that solve the index recommendation problem start with collecting a set of index candidates to select from. Each candidate within this set usually stands for a table column accessed by any of the queries in the workload.
Selection Constraints: Index selection in databases is subject to various practical constraints, such as index numbers and storage budget. Selection constraints in the paper is referred as the additional storage overhead (i.e., disk space) unless otherwise specified. During index selection, it is significant to balance performance gains against the increased resource usage, which necessitates prioritizing indexes that offer the highest benefit-to-cost ratio.
@@ 2.2. Index Advisor
Index advisors receive the workload of a specified database along with selection constraints, and then automatically generate a set of recommended indexes. The functionality of an index advisor can be formally defined as follows:
@@
(Index Advisor): Given a database , a workload consisting of a set of queries, and a storage budget , the index advisor problem seeks to automatically identify an optimal set of indexes that minimizes workload execution cost while respecting storage constraints. Formally, let denote the set of all possible column combinations that can be indexed in database . The index advisor solves the following optimization problem:
where represents the power set of all possible column combinations, computes the total execution time of workload on database with index set , and calculates the total storage cost to construct the indexes selected.
@@ 3. Related Work
3.1. Heuristic-based Index Advisors
@@
Heuristic-based advisors iteratively select candidate indexes for an input workload in a greedy fashion. Within each optimization step, the index advisors will update the index recommendation according to the estimated benefits of the index candidates. There are mainly two heuristics for the index advisors to follow: ADD heuristic and DROP heuristic. As an example of ADD heuristic, Extend (Schlosser et al., 2019) starts with no indexes and add one index at a time according to the highest benefit-to-cost ratio until the budget constraint is met. In contrast, Drop (Choenni et al., 1993) following the DROP heuristic starts with all index candidates and iteratively removes the least beneficial one until only the desired number of indexes are left. In general, heuristic-based advisors generally take significantly longer computation time to recommend indexes, especially for databases with complicated schemas. In addition, heuristic-based methods rely solely on database optimizers to evaluate index benefits, and wrong estimations in column cardinality may mislead index advisors to select indexes that are not beneficial to the workload (Ding et al., 2019).
3.2. Learning-based Index Advisors
State-of-the-art learning-based advisors (Kossmann et al., 2022; Wang et al., 2024b; Lan et al., 2020; Sadri et al., 2020a; Paludo Licks et al., 2020; Wu et al., 2022; Wang and Yoneki, 2024) predominantly adopt reinforcement learning (RL) algorithms for lower inference latency. For example, DQN (Lan et al., 2020) integrates Q-learning with deep neural networks to derive optimal policies from workload representations, while SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b) leverage the PPO (Schulman et al., 2017) algorithm for policy optimization, with BALANCE further employing a transfer learning framework to exploit historical workloads. However, RL-based methods need significantly longer training time and have lower generalisability, as training is required for each unseen database. Alternative learned models are proposed to address the limitations. For example, DBA Bandits (Perera et al., 2021b) employs a Multi-Armed Bandits framework, where index selection policies are defined and refined through neural networks to enhance computational efficiency. AutoIndex (Zhou et al., 2022) leverages a Monte Carlo Tree Search algorithm, adopting an ADD-style strategy similar to heuristic-based methods for index selection. More recently, MFIX (Chang et al., 2024) introduces a compact tree-structured index space and improves search efficiency by utilizing a multi-fidelity Bayesian optimization (MFBO) algorithm. Despite these advances, such methods often suffer from poor generalization or trade-offs between efficiency and effectiveness. Moreover, they still require retraining when the underlying data changes and fail to operate in zero-shot scenarios. It is noteworthy that some recent work focuses on specific sub-tasks of index recommendation including workload compressing (Siddiqui et al., 2022a; Brucato et al., 2024), index candidate filtering (Siddiqui et al., 2022b) and index evaluation (Zhao et al., 2022). As these studies require considerable extra design to establish a complete index advisor, we leave them for future discussion.
3.3. Prompt-based Index Advisors
Recently, LLMIdxAdvis (Zhao et al., 2025) first utilizes LLMs to generate index suggestions. Specifically, it leverages a prepared demonstration pool and employs cosine similarity to select the most relevant in-context demonstrations for a given workload. While this strategy improves efficiency and generalizability in index selection, these gains come at the expense of recommendation quality. Specifically, LLMIdxAdvis struggles to handle complex workloads and database schemas due to its reliance on limited demonstrations. In addition, it progressively builds index in a single-agent, greedy manner without refinement or verification, leading to significant hallucination problems. Deviating from LLMIdxAdvis, our work explores a multi-agent framework, which doesn’t rely on demonstrations and mitigates hallucinations via cross-agent coordination.
4. Framework Overview
In this section, we will give an overview of how we design our AMAZe framework to address the current challenges for index recommendation problem. In Section 4.1, we will first explain how we solve the zero-shot generalization challenge using a multi-agent pipeline. In Section 4.2 we will then introduce our novel framework designs, namely the workload representation paradigm and effectiveness-aware agents for better effectiveness and efficiency.